The perimeter of a college athletic field is 100 meters and the length is 14 m more than the width. Find the length and width.
What is the​ width?
What is the​ length?

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

2w+2(w+14)=100
2w+2w+28=100
4x=100-28
x=18 width
x+14=32 length

Answer by stanbon(75887)

The perimeter of a college athletic field is 100 meters and the length is 14 m more than the width. Find the length and width.
-------------------------------------
P = 2(L+W)
100 = 2(W+14+W)
2W + 14 = 50
2W = 36
Width = 18 m
---
Length = W+14 = 32 m
------
Cheers,
Stan H.
--------------

What is the​ width?
What is the​ length?
------